OpenStudy (anonymous):

derivative of cos^2(x^3)

OpenStudy (anonymous):

Use the chain rule.

OpenStudy (anonymous):

i understand to do that but i get stuck

OpenStudy (anonymous):

i got -2sin^2(x^3)(3x^2) -6x^2sin^2(x^3)

OpenStudy (anonymous):

(-6x^2)sin^2(x^3)

OpenStudy (anonymous):

You should have a term that still has \(\cos(x^3)\) in it. Remember the chain rule: \((f\circ g)'(x)=f'(g(x))g'(x)\). If you need to, write down explicitly at each step what f, g, f', and g' are.

OpenStudy (anonymous):

The derivative of \(\cos^2x\) is not \(-\sin^2x\). You have to use the chain rule.

OpenStudy (anonymous):

thats what im asking for help with.